OpenStudy (anonymous):

Why do you think 2 plus 2 equals two? Write at least 3 sentences. (this question is seriously on here)

OpenStudy (aroub):

1. if u add 2 to 2 u will get 4 what so ever (for example hold up 2 fingers and 2 fingers what would u get? )

OpenStudy (anonymous):

but why?

OpenStudy (anonymous):

2+0+2+0=4

OpenStudy (aroub):

oh good super!

OpenStudy (osanseviero):

1. There is a demonstrated way that 1 + 1 can be 3 http://openstudy.com/groups/mathematics#/groups/mathematics/updates/4e02adc70b8bbe4f12e8bc71 I think there is more than one sentence in the link :)

OpenStudy (anonymous):

I don't think so. I don't think so. I don't think so.

OpenStudy (anonymous):

yea u can. it says y or y not you agree

OpenStudy (anonymous):

or dont agree

OpenStudy (anonymous):

Within computer science if you set the term "two" to equal the numerical value of "4" then 2+2=4 and 4="two"

OpenStudy (osanseviero):

It is not exact... 2 quarters of water plus 2 quarters of water is never 4 quarters of water... It is matter in motion. Some can evaporate... A little will move to another place, etc

OpenStudy (anonymous):

i can also prove 2+2=undefined

OpenStudy (anonymous):

is undefined = two ?

OpenStudy (anonymous):

you must use at least 3 sentences in the proof.

OpenStudy (anonymous):

proof 2+2 4 now divide it by 0 4/0=undefined(anything divided by 0 is undefined)